### v1

@@ Abstract
Accurate and high-resolution Earth system model (ESM) simulations are essential to assess the ecological and socio-economic impacts of anthropogenic climate change, but are computationally too expensive. Recent machine learning approaches have shown promising results in downscaling ESM simulations, outperforming state-of-the-art statistical approaches.
However, existing methods require computationally costly retraining for each ESM and extrapolate poorly to climates unseen during training. We address these shortcomings by learning a consistency model (CM) that efficiently and accurately downscales arbitrary ESM simulations without retraining in a zero-shot manner.
Our foundation model approach yields probabilistic downscaled fields at resolution only limited by the observational reference data. We show that the CM outperforms state-of-the-art diffusion models at a fraction of computational cost while maintaining high controllability on the downscaling task. Further, our method generalizes to climate states unseen during training without explicitly formulated physical constraints.
@@ 1 Introduction
Accurate and high-resolution climate simulations are of crucial importance to project the climatic, hydrological, ecological, and socioeconomic impacts of anthropogenic climate change. Precipitation, in particular, is one of the most important climate variables, with huge impacts, e.g., on vegetation and crop yields, infrastructure, or the economy [Kotz et al., 2022]. However, it is also the variable that is arguably most difficult to model and predict, especially extreme precipitation events. Numerical Earth system models (ESMs) are our main tool to project the future evolution of precipitation and its extremes. However, they exhibit biases and have much coarser spatial resolution, on the order of 10-100km, than needed for reliable assessments of the impacts of climate change [Schneider et al., 2017].
@@
Due to the chaotic nature of geophysical fluid dynamics, the trajectory of climate simulations will not match historical observations, requiring approaches suitable for unpaired samples to learn such tasks. Recently, methods from generative deep learning have shown promising results in downscaling or correcting spatial patterns in ESM simulations. Normalizing flows (NFs) [Dinh et al., 2015] and generative adversarial networks (GANs) [Goodfellow et al., 2014] can perform these tasks efficiently in a single step [Groenke et al., 2020, Pan et al., 2021, François et al., 2021, Harris et al., 2022, Hess et al., 2022, Hess et al., 2023]. However, NFs often exhibit lower quality in the generated images, while GANs can suffer from training instabilities and problems such as mode collapse [Arjovsky and Bottou, 2017]. Moreover, these approaches require computationally expensive retraining of the neural networks for each ESM to be processed. This makes downscaling large ESM ensembles, as needed in impact assessments, prohibitively costly and time-consuming.
@@
Diffusion-based generative models have demonstrated superior performance over NFs and GANs on classical image generation tasks [Dhariwal and Nichol, 2021, Song et al., 2021]. Crucially, iteratively solving the reversed diffusion equation allows for strong control over the image sampling process. This enables so-called foundation models that are only trained on a given target dataset and which can later be repurposed, e.g., to generate realistic images based on a given “stroke sketch” guide as in SDEdit [Meng et al., 2022].
@@
So far, such foundation model-based approaches have only been applied to downscale idealized fluid dynamics [Bischoff and Deck, 2023, Wan et al., 2023] in Earth system science-related tasks. Both studies use stochastic differential equation (SDE) based diffusion models and achieve remarkable performance. The iterative integration of the SDE, however, implies that the generative network needs to be evaluated up to 1000 times in order to downscale a single simulated field. This makes such methods unsuitable for processing large simulation datasets, e.g., at high temporal resolution or over long periods, as would be needed in the context of climate change projections and impact assessments.
In this work, we tackle these shortcomings using a novel type of generative approach based on consistency models (CM) [Song et al., 2023] to downscale global precipitation simulations from a fully-coupled ESM in a single step, without sacrifices in performance or the controllability of the sampling. We train the CM network from scratch (“in isolation”) on ERA5 reanalysis data only, making the training independent of any ESM. In summary, our study makes the following contributions:
@@
  • Our foundation model is trained on the target dataset only, without conditioning on the ESM. This makes our method applicable to any ESM without requiring computationally expensive retraining.
  • The training is much more stable than in previous methods based on CycleGAN [Zhu et al., 2017, Pan et al., 2021, François et al., 2021, Hess et al., 2022, Hess et al., 2023]).
@@
  • The generative downscaling is controllable at the inference stage after training, so that the downscaled fields preserve spatial patterns of the ESM up to a chosen characteristic spatial scale.
  • The consistency model employed here is up to three orders of magnitude faster than current state-of-the-art diffusion model-based approaches [Bischoff and Deck, 2023], yet shows superior performance.
@@
  • The highly efficient CM allows the generation of a large number of downscaled realizations for a single ESM field (i.e., a one-to-many mapping), enabling a robust quantification of the sampling spread.
  • Our method is robust to out-of-sample predictions, showing strong preservation of the non-stationary dynamics in future transient climates (e.g., in an extreme SSP5-8.5 scenario). This is a big advantage over previous deep learning-based methods that require specifically formulated physical constraints in order to generalize [Beucler et al., 2021, Harder et al., 2022, Hess et al., 2022].
@@
Figure 1: Sketch of the foundation model for downscaling of Earth system model fields. (Upper panel) Unconditional training of the score-based diffusion and consistency models (CM) that learn to reverse a forward diffusion process. While the stochastic differential equation of the diffusion model requires an iterative integration over many steps, the CM only takes a single step to generate a global precipitation field from noise. (Lower panel) The unconditionally trained foundation model is used to downscale (upsample) a low-resolution ESM precipitation field to a four times higher resolution. By adding noise of a chosen variance to the ESM field, the spatial scale to be preserved in the ESM can be controlled: small noise variance implies a close pairing to the original ESM field with only small changes; a larger variance will result in changes of larger spatial scales and in weaker pairing to the ESM field.
@@ 2 Results
We evaluate the efficient consistency model-based downscaling method against the SDE bridge from [Bischoff and Deck, 2023] over the test set data. We investigate the performance of the downscaling, the ability to correct distributional biases in the ESM, the sample spread given a chosen spatial scale, and the preservation of trends in future climate scenarios.
@@ 2.1 Downscaling Spatial Fields
For a qualitative comparison, we show single precipitation fields in Fig. 2. Both generative downscaling methods based on the SDE bridge (Fig. 2E) or CM (Fig. 2G) are able to produce high-resolution precipitation fields that are visually indistinguishable from the unpaired ERA5 field (Fig. 2A).
When upscaled back to the native POEM resolution using a $4\times 4$-kernel average pooling, an accurate representation of the low-resolution POEM simulation field is apparent. As indicated in Fig. 2F and Fig. 2H, a high Pearson correlation of $0.89$ and $0.95$ for the SDE and CM methods, respectively, is maintained between upscaled corrected and native fields.
@@
We provide correlation statistics for the entire test set in Table 1. Besides the average pooled fields, we also compare the downscaled and linearly interpolated POEM simulation on the high-resolution grid by applying a low-pass filter with a cut-off frequency set to $k^{*}=0.0667$ on the downscaled fields before computing the correlation. In this way we compare the preservation of the large-scale patterns in the ESM. The SDE bridge achieves a mean correlation of $0.918$ and $0.916$ for the average pooled and low-pass filtered fields, respectively. Our CM-based method achieves even higher correlation values of $0.954$ and $0.941$ for both measures.
@@
We estimate the average time it takes to produce a single sample with the SDE and CM methods on a NVIDIA V100 32GB GPU. The average is taken over $100$ samples and we set the number of SDE integration steps to $500$ as in [Bischoff and Deck, 2023], which is lower then the typical $1000-2000$ steps. The SDE takes on average $39.355$ seconds, while the CM samples much more efficiently taking only $0.116$ seconds.
@@
We analyze the downscaling performance of the CM and SDE bridge approaches quantitatively using power spectral densities (PSDs) as in [Ravuri et al., 2021, Hess et al., 2023]. The ESM fields under-represent variability at small spatial scales. This implies an underestimation of spatial intermittency, i.e., overly smooth precipitation patterns. This issue is well known and presents a key problem for the assessment of the impacts of extreme precipitation in a changing climate. The generative downscaling methods based on SDEs and CM perform very well in increasing spatial resolution and greatly improve the spatial intermittency at the smaller spatial scales (Fig. 3A).
@@
We also investigate the change in PSDs as a function of the noise variance schedule time $t$ in Fig. 3B, which is associated with the spatial scale up to which patterns are corrected, as explained in the Methods section. For minimal noise, i.e., $t^{*}=t_{\mathrm{min}}$, the CM model reproduces the PSD of the ESM as one would expect, as there are no changes made to the ESM field. For maximal variance with $t^{*}=t_{\mathrm{max}}$, the PSD closely matches the ERA5 reanalysis ground truth. For any $t_{\mathrm{min}}<t^{*}<t_{\mathrm{max}}$, we find a trade-off between the two extreme cases that match the PSD above and below the intersection (dotted grey line) to a certain degree, depending on the spatial scale preserved in the ESM.
@@
Figure 2: Qualitative comparison of single-day precipitation fields. (A) Daily precipitation from the ERA5 target datasets was used for training the generative foundation models, and (B) at four times lower resolution for comparisons. (C) A precipitation field from a historical run of the POEM ESM interpolated to the target resolution and (D) on its native resolution of $3^{\circ}\times 3.75^{\circ}$. The POEM fields is unpaired with the ERA5 field from the same date or any other ERA5 field. (E) Downscaled field from POEM (D) with the SDE bridge method. (F) An upscaled (average pooled) representation of (E) for comparison with the original POEM field shown in D and the Pearson correlation between the two. In (G) and (H), the respective fields for the CM method are shown. Note that the CM downscaling yields a higher correlation, and hence better consistency of the large-scale features, than the SDE method.
[…]
SDE 0.918 0.916 0.214 72.51 1.106 68.15 39.355
CM 0.954 0.941 0.217 72.08 1.080 68.92 0.116
@@
Figure 3: Mean spatial power spectral densities (PSDs) of global precipitation fields. (A) Comparison of the PSDs for the target ERA5 reanalysis data (black), the POEM simulations interpolated to the same high-resolution grid (orange), the SDE bridge (cyan), and the CM downscaling (magenta). The vertical dashed lines marks the spatial scale at which the PSDs of POEM and ERA5 intersect and is thus a natural choice for for the wavenumber $k^{\ast}$ up to which to correct, which in turn determines $t^{\ast}$, i.e. the noising strength in the diffusion models (see Eq. 8 in the Methods). (B) CM downscaling (dashed lines) applied to be consistent with different spatial scales as a function of the noising strength $t$ over the entire range $[t_{\mathrm{min}},t_{\mathrm{max}}]$. Noising small scales implies nearly reproducing the POEM simulations while noising larger scales corresponds to a weaker pairing to the ESM (see Fig. 1).
@@ 2.2 Bias Correction
We compare the ability to correct biases in the ESM with histograms of relative frequency and latitude-profiles of mean precipitation to investigate the reduction of known biases such as a double-ITCZ [Tian and Dong, 2020], following the evaluation methodology in [Hess et al., 2022, Hess et al., 2023].
For the reasons explained above, when applied to the ESM simulations without QDM preprocessing, the ability of the CM to correct biases naturally depends on the chosen spatial consistency scale (Fig. S2). Selecting the smallest scale reproduces the ESM without any changes, hence inheriting its biases. Choosing the largest possible scales generates samples with statistics very close to the target dataset. However, the fields become more and more unpaired to the ESM at such high noise levels (see Fig. 1). A scale between these two extremes will correct for biases to a varying degree, depending on the chosen correction scale.
@@
Applying QDM to POEM strongly improves the frequency distributions as expected. Downscaling the ESM with the SDE further improves the global histograms by an order magnitude, particularly for the extremes. Our CM-based method shows the overall largest bias reduction in the global histograms (Fig. 4B).
[…]
The ESM shows a strong double-ITCZ that is common among state-of-the-art ESMs [Tian and Dong, 2020] (Fig.4C). As expected, QDM is able to remove most of the biases, though slightly underestimating the peak north of the equator in the ITCZ. The downscaling methods based on the SDE bridge and CM show a similar absolute error for these latitude profiles as when only applying QDM alone (Fig.4D).
We report the absolute value of the grid cell-wise error in the long-term mean, again averaged globally, for both downscaling methods (Tab. 1). The SDE downscaling results in an error of $0.214$, reducing the error in the POEM ESM by $72.51\%$, and performing slightly better than our CM method which exhibits an error of $0.217$ and a respective error reduction of $72.08\%$ in the POEM ESM.

### v3

@@ Abstract
Accurate and high-resolution Earth system model (ESM) simulations are essential to assess the ecological and socio-economic impacts of anthropogenic climate change, but are computationally too expensive to be run at sufficiently high spatial resolution. Recent machine learning approaches have shown promising results in downscaling ESM simulations, outperforming state-of-the-art statistical approaches. However, existing methods require computationally costly retraining for each ESM and extrapolate poorly to climates unseen during training. We address these shortcomings by learning a consistency model (CM) that efficiently and accurately downscales arbitrary ESM simulations without retraining in a zero-shot manner. Our approach yields probabilistic downscaled fields at a resolution only limited by the observational reference data. We show that the CM outperforms state-of-the-art diffusion models at a fraction of computational cost while maintaining high controllability on the downscaling task. Further, our method generalizes to climate states unseen during training without explicitly formulated physical constraints.
Contact: Philipp Hess, philipp.hess@tum.de
@@ 1 Introduction
Accurate and high-resolution climate simulations are of crucial importance in projecting the climatic, hydrological, ecological, and socioeconomic impacts of anthropogenic climate change. Precipitation is one of the most important climate variables, with particularly large impacts, e.g., on vegetation and crop yields, infrastructure, or the economy [kotz_effect_2022]. However, it is also the variable that is arguably most difficult to model and predict, especially extreme precipitation events, resulting from a complex interaction of processes over a large range of spatial and temporal scales that cannot be fully resolved.
@@
Numerical Earth system models (ESMs) are our main tool to project the future evolution of precipitation and its extremes. However, they exhibit biases and have much coarser spatial resolution, on the order of 10-100km, than needed for reliable assessments of the impacts of climate change [schneider_climate_2017]. Therefore, ESM projections have to be downscaled to higher resolution, for which several machine learning-based approaches have recently been proposed. Due to the chaotic nature of geophysical fluid dynamics, the trajectory of climate simulations will not match historical observations, requiring approaches suitable for unpaired samples to learn such tasks. Recently, methods from generative deep learning have shown promising results in downscaling or correcting spatial patterns of ESM simulations. Normalizing flows (NFs) [dinh_nice_2015] and generative adversarial networks (GANs) [goodfellow_generative_2014] can perform these tasks efficiently in a single step [groenke_climalign_2020, pan_learning_2021, francois_adjusting_2021, harris_generative_2022, hess_physically_2022, hess_deep_2023]. However, NFs often exhibit lower quality, e.g., less sharp and detailed generated images, while GANs can suffer from training instabilities and problems such as mode collapse [arjovsky_towards_2017]. Moreover, these approaches require computationally expensive retraining of the neural networks for each specific ESM simulation to be processed. This makes downscaling large ESM ensembles, as needed in impact assessments, prohibitively costly and time-consuming.
@@
Diffusion-based generative models have demonstrated superior performance over NFs and GANs on classical image generation tasks [dhariwal_diffusion_2021, song_score-based_2021]. Crucially, iteratively solving the reversed diffusion equation allows for strong control over the image sampling process.
This enables foundation models for computer vision and image processing, which are only trained to generate target dataset samples from noise and which can later be repurposed for different downstream tasks, e.g., inpainting, classification, colorization, or generating realistic images based on a given “stroke sketch” guide as with SDEdit [clark_text--image_2023, sauer_adversarial_2024, meng_sdedit_2022].
@@
So far, such diffusion model-based approaches have only been applied to downscale idealized fluid dynamics [bischoff_unpaired_2024, wan_debias_2023] in Earth system science-related tasks. Both studies use stochastic differential equation (SDE) based diffusion models and achieve remarkable performance. The iterative integration of the SDE, however, implies that the generative network needs to be evaluated up to 1000 times in order to downscale a single simulated field. This makes such methods unsuitable for processing large simulation datasets, e.g., at high temporal resolution or over long periods, as would be needed in the context of climate change projections and impact assessments.
Much effort has been taken to improve the sampling speed of diffusion models. Using ordinary differential equations instead of SDEs can reduce the number of integration steps to around 10-50 [karras_elucidating_2022, esser_scaling_2024]. Distillation techniques can also be applied to improve the sampling speed to a single step [luhman_knowledge_2021, zheng_fast_2024]. However, they require additional training costs and hyperparameter choices. In general, a trade-off between the number of generative steps and the resulting image quality scores has been found [song_score-based_2021, esser_scaling_2024]. Recently, consistency models (CMs) have been proposed that can be directly trained as single-step generative models without sacrifices in the controllability of the sampling [song_consistency_2023]. CMs have been shown to outperform distilled single-step diffusion models on common image benchmarks [song_consistency_2023].
In this work, we tackle the shortcomings of standard diffusion models by using CMs [song_consistency_2023] to downscale global precipitation simulations from three different ESMs in a single step. In summary, we aim to establish a generative model for downscaling ESM fields that fulfills the following conditions:
@@
  • The CM model is trained on the target dataset only, without conditioning on the ESM. This makes the method applicable to any ESM without requiring computationally expensive retraining.
  • The training minimizes a regression loss that is more stable than the adversarial training in previous methods [pan_learning_2021, francois_adjusting_2021, hess_physically_2022, hess_deep_2023].
@@
  • The generative downscaling is controllable at the inference stage after training, allowing the choice of a characteristic spatial scale up to which unbiased spatial patterns are preserved in the ESM.
  • The model only requires a single network evaluation instead of the many iterative evaluations in diffusion model-based downscaling approaches [bischoff_unpaired_2024, wan_debias_2023].
@@
  • The model allows to generate a large number of downscaled realizations for a single ESM field (i.e., a one-to-many mapping), enabling probabilistic downscaling and a quantification of the sampling spread.
  • The method does not require specifically formulated physical constraints [harder_hard-constrained_2023] in order to preserve trends [hess_physically_2022].
@@
Figure 1: Sketch of the consistency model for downscaling of Earth system model fields. (Upper panel) Unconditional training of the score-based diffusion and consistency models (CM) that learn to reverse a forward diffusion process. While the stochastic differential equation of the diffusion model requires an iterative integration over many steps, the CM only takes a single step to generate a global precipitation field from noise. (Lower panel) The unconditionally trained consistency model is used to downscale (upsample) a low-resolution ESM precipitation field to a four times higher resolution. By adding noise of a chosen variance to the ESM field, the spatial scale to be preserved in the ESM can be controlled: small noise variance implies a close pairing to the original ESM field with only small changes; a larger variance will result in changes at larger spatial scales and in weaker pairing to the ESM field.
@@ 2 Results
We evaluate our CM-based downscaling method against the SDE bridge from [bischoff_unpaired_2024] over the test set data. We investigate the performance of the downscaling, the ability to correct distributional biases in the ESM, the sample spread given a chosen spatial scale, and the preservation of trends in future climate scenarios.
@@ 2.1 Downscaling Spatial Fields
For a qualitative comparison, we show single precipitation fields in Fig. 2. Both generative downscaling methods, based on the SDE bridge (Fig. 2E) and CM (Fig. 2G), respectively, are able to produce high-resolution precipitation fields that are visually indistinguishable from the unpaired ERA5 field (Fig. 2A). A similar performance can be seen when downscaling simulations from the state-of-the-art GFDL-ESM4, the more lightweight SpeedyWeather.jl general circulation model, as well as when applying our method to initially upscaled ERA5 data as a proof of concept (Fig. S3, S6).
@@
When upscaled back to the native ESM resolution using a $4\times 4$-kernel average pooling, an accurate representation of the low-resolution ESM simulation field is apparent. As indicated in Fig. 2F and Fig. 2H for the case of the POEM model, a high Pearson correlation of $0.89$ and $0.95$ for the SDE and CM methods, respectively, is maintained between upscaled corrected and native fields. We provide correlation statistics for the entire test set in Table LABEL:tab:correlation. Besides the average pooled fields, we also compare the downscaled and linearly interpolated POEM simulation on the high-resolution grid by applying a low-pass filter with a cut-off frequency set to $k^{*}=0.0667$ on the downscaled fields before computing the correlation. In this way, we test the preservation of the large-scale patterns in the ESM. More specifically, we measure how similar the downscaled fields are to the original coarse fields when upscaling them back to the same resolution. The SDE bridge achieves a mean correlation of $0.918$ and $0.916$ for the average pooled and low-pass filtered fields, respectively. Our CM-based method achieves even higher correlation values of $0.954$ and $0.941$ for both measures. The CM-based downscaling also shows higher correlations when downscaling ERA5 (see Table S2).
@@
We estimate the average time it takes to produce a single sample with the SDE and CM methods on a NVIDIA V100 32GB GPU. The average is taken over $100$ samples, and we set the number of SDE integration steps to $500$ as in [bischoff_unpaired_2024], which is lower than the typical $1000-2000$ steps [dhariwal_diffusion_2021, song_score-based_2021]. The SDE takes on average $39.4$ seconds, while the CM samples much more efficiently, taking only $0.1$ seconds, scaling linearly with the number of samples.
@@
We analyze the downscaling performance of the CM and SDE bridge approaches quantitatively using power spectral densities (PSDs) as e.g. in [ravuri_skilful_2021, hess_deep_2023]. The interpolated ESM fields under-represent variability at small spatial scales, since these are not present in the low-resolution simulation. This implies an underestimation of spatial intermittency, i.e., overly smooth precipitation patterns, which is highly problematic for impact assessment. The generative downscaling methods based on SDEs and CM perform very well in increasing spatial resolution and greatly improve the spatial intermittency at the smaller spatial scales (Fig. 3A). The PSD improvements are found to be consistent when applying the CM downscaling to the GFDL-ESM4 and SpeedyWeather.jl simulations (Fig. S4).
@@
We also investigate the change in PSDs as a function of the noise variance schedule time $t$ in Fig. 3B, which is directly related the spatial scale up to which patterns are corrected (see Methods). For minimal noise, i.e., $t^{*}=t_{\mathrm{min}}$, the CM model reproduces the PSD of the ESM, as expected from Eq. 5 since there are no changes made to the ESM field. For maximal variance with $t^{*}=t_{\mathrm{max}}$, the PSD closely matches the ERA5 reanalysis ground truth. For any $t_{\mathrm{min}}<t^{*}<t_{\mathrm{max}}$, we find a trade-off between the two extreme cases that match the PSD above and below the intersection (dotted grey line) to a certain degree, depending on the spatial scale preserved in the ESM.
To test the temporal consistency of the downscaled fields, we report temporal autocorrelation values computed for each grid cell as global averages in Fig. S8. We find that the generative and stochastic SDE and CM methods result in a more accurate temporal correlation than the deterministic approach of quantile mapping and bilinear interpolation.
@@
Figure 2: Qualitative comparison of single-day precipitation fields. (A) Daily precipitation from the ERA5 target dataset was used for training the generative models. (B) Same as (A) but at four times lower resolution for comparisons. (C) A precipitation field from a historical run of the POEM ESM interpolated to the target resolution and (D) on its native resolution of $3^{\circ}\times 3.75^{\circ}$. The POEM fields are unpaired with the ERA5 field from the same date or any other ERA5 field. Downscaled field from POEM (D) with the SDE bridge method (D)$\rightarrow$(E). (F) An upscaled (average pooled) representation of (E) for comparison with the original POEM field is shown in D and the Pearson correlation between the two. Downscaling POEM with the CM-based method (D)$\rightarrow$(G), and the respective pooled field (H). Note that the CM downscaling yields a higher correlation, and hence better consistency of the large-scale features, than the SDE method.
Table 1: Summary statistics comparing our CM approach and the SDE bridge as a benchmark. Correlations are computed (2nd column) on the native POEM grid ($r=3^{\circ}\times 3.75^{\circ})$ by applying a $4\times 4$ average pooling to the downscaled fields and (3rd column) on the downscaled grid by applying a low-pass filter to the high-resolution downscaled fields with a cut-off frequency that is consistent with the chosen spatial scale in POEM to be preserved. (4th column) The global long-term mean absolute error with respect to ERA5 and (5th column) the respective error reduction compared to the POEM ESM bias. (6th column) The global absolute error in the 95th precipitation percentile and (7th column) the error reduction with respect to the POEM ESM bias. An estimation of the mean sampling time is also reported (8th column).
 [^1] The best performance is given in bold font.
Model Correlation (pooled) Correlation (low-pass) Mean error % 95th precentile error % Sample time [s]
SDE 0.918 0.916 0.214 72.51 1.106 68.15 39.4
CM 0.954 0.941 0.217 72.08 1.080 68.92 0.1
@@
Figure 3: Mean spatial power spectral densities (PSDs) of global precipitation fields. (A) Comparison of the PSDs for the target ERA5 reanalysis data (black), the POEM simulations interpolated to the same high-resolution grid (orange), the SDE bridge (cyan), and the CM downscaling (magenta). The vertical dashed lines mark the spatial scale at which the PSDs of POEM and ERA5 intersect and are thus a natural choice for the wavenumber $k^{\ast}$ up to which to correct, which in turn determines $t^{\ast}$, i.e. the noising strength in the diffusion models (see Eq. 11 in the Methods). (B) CM downscaling (dashed lines) applied to be consistent with different spatial scales as a function of the noising strength $t$ over the entire range $[t_{\mathrm{min}},t_{\mathrm{max}}]$. Noising small scales implies nearly reproducing the POEM simulations, while noising larger scales corresponds to a weaker pairing to the ESM (see Fig. 1).
@@ 2.2 Bias Correction
We compare the ability to correct biases in the ESM with histograms of relative frequency and latitude-profiles of mean precipitation to investigate the reduction of known biases such as a double-ITCZ [tian_double-itcz_2020], following the evaluation methodology in [hess_physically_2022, hess_deep_2023].
When applied to the ESM simulations without preprocessing via quantile delta mapping (QDM) [cannon_bias_2015] , the ability of the CM to correct biases naturally depends on the chosen spatial consistency scale (Fig. S9). Selecting the smallest scale reproduces the ESM without any changes, hence inheriting its biases. Choosing the largest possible scales generates samples with statistics very close to the target dataset. However, the fields become more and more unpaired to the ESM at such high noise levels (see Fig. 1). A scale between these two extremes will correct for biases to a varying degree, depending on the chosen correction scale.
@@
Applying QDM to POEM strongly improves the frequency distributions as expected. Downscaling the ESM with the SDE further improves the global histograms by an order magnitude, particularly for the extremes. Our CM-based method shows the overall largest bias reduction in the global histograms (Fig. 4B). When applied to the different ESMs and the upscaled ERA5 data, our method shows a consistent bias correction skill (Fig. S5).
We further compute the error in the 95th percentile of the local precipitation histogram for each grid cell and aggregate the absolute value globally (Tab. LABEL:tab:correlation). The SDE method shows an error of $1.106$mm/day, reducing the error of the POEM ESM by $68.15\%$. The CM method performs slightly better with an error of $1.08$mm/day and a respective error reduction of $68.92\%$ of the ESM. When downscaling the initially upsampled ERA5 data, the CM again shows better performance, with an error of $0.725$mm/day compared to the SDE with an error of 0.868mm/day (Tab. S2).
The POEM model exhibits a strong double-ITCZ bias that is common among state-of-the-art ESMs [tian_double-itcz_2020] (Fig.4C). As expected, QDM is able to remove most of the biases, though slightly underestimating the peak north of the equator in the ITCZ. The downscaling methods based on the SDE bridge and CM show a similar absolute error for these latitude profiles as when only applying QDM alone (Fig.4D).
We report the absolute value of the grid cell-wise error in the long-term mean, again averaged globally, for both downscaling methods (Tab. LABEL:tab:correlation). The SDE downscaling results in an error of $0.214$mm/day, reducing the error in the POEM model by $72.51\%$, and performing slightly better than our CM method, which exhibits an error of $0.217$,,/day and a respective error reduction of $72.08\%$ when applied to POEM. Evaluated on the ERA5 data, we find that the CM method performs slightly better than the SDE in terms of mean absolute error and RMSE (Tab. S2).